OpenStudy (anonymous):

Find the vertical asymptotes of the rational function below:

OpenStudy (anonymous):

\[g(x)=\frac{ -5x }{ x^{2}-x-6}\]

OpenStudy (anonymous):

Vertical asymptotes is x values when the denominator = 0

OpenStudy (anonymous):

Yeah but I can't figure out what they are because I'm retarded.

OpenStudy (anonymous):

\[x ^{2}-x-6=0\]

OpenStudy (anonymous):

\[x ^{2}-x=6\]

OpenStudy (anonymous):

\[\sqrt{x ^{^{2}}-x}=\sqrt{6}\]

OpenStudy (anonymous):

x² -x - 6 = 0 ( x - 3) ( x + 2) = 0

OpenStudy (anonymous):

\[x-\sqrt{x}=\sqrt{6}\]

OpenStudy (anonymous):

>.<!

OpenStudy (anonymous):

Since the product C = - 6 Sum B = -1 => x = 3, x = -2

OpenStudy (anonymous):

@natnatwebb It's necessary to learn factorize by product and Sum method!

OpenStudy (anonymous):

And that, wonderful, wonderful OpenStudy member, was what I was missing. THANK YOU SO MUCH! >.<!